OpenStudy (anonymous):

1 + cos θ - 2sin^(2)θ=0

OpenStudy (abb0t):

I would change\[\sin^2(\theta)\] to one of the cosine trig identities.

OpenStudy (anonymous):

Which one would that be? Sin^2 x + Cos^2 x =1? Im confused lol

OpenStudy (abb0t):

\[\sin^2(\theta) = 1-\cos^2(x)\] then you'll have it in the form: \[ax^2+bx+c\] and you can factor it out, solve when it's equal to zero.

OpenStudy (anonymous):

Alright I see what you did there, let me check it out

OpenStudy (anonymous):

Got it, thank you